A particle of 0.5 kg is thrown horizontally with speed of 10m/s in the vertical plane. Point of projection is taken as origin and j is vertically upward. Angular momentum of the particle after 1 s about origin is

Grade:11

1 Answers

Vikas TU
7 years ago
Distance covered by the particle in 1 second in x and y direction respectively is:
x = 10*1 = 10 meter.
and
y = 0 + 0.5*g*1^2 = > 5 meter.
net displacement => root(100 + 25) = > root(125) => 5*root(5)
The angular momentum avout the origin is:
=> mvr 
=> 0.5*10*5*root(5)
=> 25root(5) Kg/s